An ideal heat engine absorbs 84.5 kJ of heat and exhausts 71.8 kJ of heat in each cycle. What is the efficiency of the engine?

Tries 0/12

How much work is done in a cycle?
Answer #1

(a)

formula for efficiency is

e = 1- QC/QH

= 1- 71.8/84.5

= 0.15 or 15 %

(b)

W = e QH = 0.15029 (84.5 kJ) = 12.7 kJ